OpenStudy (volleyballlover55):

If y − 1 = 4x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(1, 4), (2, 8), (3, 12)} {(4, 1), (8, 2), (12, 3)} {(0, 1), (1, 5), (2, 9)} {(1, 0), (5, 1), (9, 2)}

OpenStudy (volleyballlover55):

@Michele_Laino

OpenStudy (michele_laino):

hint: please compute these quantities: \[\begin{gathered} y = 4 \cdot 0 + 1 = ...? \hfill \\ y = 4 \cdot 1 + 1 = ...? \hfill \\ y = 4 \cdot 2 + 1 = ...? \hfill \\ \end{gathered} \]

OpenStudy (volleyballlover55):

1 5 9

OpenStudy (michele_laino):

so we can write this table: